40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 542     Entry time: Fri Jun 22 10:53:06 2018
Author: Alena 
Type: General 
Category: General 
Subject: Laser failure 

One of the HeNe lasers died (the one for measuring slots CR2 and CR4). It had no outgoing light when I came in this morning. I connected the laser to another identical power supply. The laser started going on and of and the power supply was making some sparkling noise. I swapped the laser with the test chamber (CR0). No re alignment was required for measurement slots CR1 and CR3. A very slight alignment was done for measuring slots CR2 and CR4 after the laser from the test chamber was installed. So temporary we do not have a setup to measure 2" samples easily. Attaching pictures of the broken laser`s label

Attachment 1: 20180622_112134.jpg  3.055 MB  Uploaded Fri Jun 22 17:26:36 2018  | Hide | Hide all
20180622_112134.jpg
Attachment 2: 20180622_112138.jpg  3.098 MB  Uploaded Fri Jun 22 17:26:36 2018  | Hide | Hide all
20180622_112138.jpg
ELOG V3.1.3-